OpenStudy (anonymous):

what is 3x3x3x3x3 in exponent form

OpenStudy (anonymous):

5 three means power of 5

OpenStudy (solomonzelman):

\[a \times a \times a \times a \times a=a^5\]

OpenStudy (anonymous):

ahh thank you

OpenStudy (solomonzelman):

\[3 \times 3 \times 3 \times 3 \times 3 = ??\]